# Monocular Lane Detection Based on Deep Learning: A Survey

### v1

@@ IV-B2 Learnable View Transformation
To reduce the inherent errors caused by this rigid transformation, some researches adopt a more flexible approach. They use neural networks to learn the transformation process from FV to BEV features. PersFormer [24] leverages deformable attention [99] to learn the spatial transformation from FV to BEV. It references the coordinate transformation matrix of IPM to generate BEV feature representations, focusing on relevant regions in the FV features. In the lane decoding stage, PersFormer adopts a unified 2D/3D line anchor design, achieving unified 2D and 3D lane detection. BEV-LaneDet [26] integrates the MLP based view transformation method VPN [120] into FPN [131] to obtain BEV features. It constructs a virtual camera module to project all images onto a standard virtual camera view, ensuring consistent image distribution. For lane modeling, BEV-LaneDet models lanes as keypoints. It adopts an embedding-clustering-based instance segmentation method [65] and refers to YOLO [96] to divide the BEV plane into grids, predicting the offset of each grid’s center point relative to GT. Due to its concise architecture, it is well-suited for deployment. Yao et al. [121] add the coarse-to-fine mechanism [19] based on PersFormer [24]. They fuse the local and global information referring to the coordinates of sparse points and jointly refine the global and local structures of lanes. Chen et al. [132] decompose the cross-attention between FV and BEV features into separate cross-attentions: one between FV and lane features, and another between BEV and lane features. Dynamic kernels are then used to convolve FV and BEV feature maps, generating 2D and 3D lane keypoint offset maps. GroupLane [133] uses the depth-estimation-based view transformation method LSS [126] to obtain BEV features. It models BEV lanes as grids [68] and establishes vertical and horizontal group heads to identify horizontal and vertical lanes, respectively. LaneCPP [125] also completes BEV transformation based on LSS. It models lanes as B-spline curves and uses prior knowledge of road geometry to enhance view transformation and lane prediction.
@@ IV-C2 Derictly Modeling 3D Lanes
The 3D object detection methods based on sparse BEV representations, such as DETR3D [137] and PETR [138], provide guidance for this approach. CurveFormer [119] constructs curve queries by modeling 3D lanes as 3D line anchors to provide explicit positional priors [139]. It designs a curve cross-attention mechanism to predict polynomial parameters for 3D lanes. Inspired by the line anchor feature pooling mechanism in LaneATT [80], Anchor3DLane [25] utilizes camera intrinsics and extrinsics to accurately project 3D line anchor points onto FV features. This facilitates anchor feature sampling, allowing the network to predict 3D coordinates and lane classification results based on the sampled anchor features. LATR [27] decomposes 3D line anchors into dynamically generated point level and lane level queries. It uses dynamic 3D ground position embeddings to interact with FV features, updating the lane query to bridge 3D space and 2D images. Dong et al. [124] model lanes as 3D Bézier curves and predict the curve control points via Transformer. Han et al. [140] expressed the 3D lane as polynomial in 3D space. They design two Transformer structures to learn the 2D polynomial and height of the X-O-Z plane, and project the resulting 3D lane onto FV for supervised alignment. PVALane [123] generates sparser 3D line anchors than Anchor3DLane by first predicting 2D lanes in FV. It also introduces a module to align sampled FV and BEV features for more accurate 3D lane detection.
@@ V Benchmark Results
This section reports the performance of representative lane detection methods on commonly used public datasets. For each reviewed area, the most widely used datasets are selected for benchmarking in Section V-A. Because of the high efficiency requirement for lane detection, speed tests are also conducted on representative open-source lane detection methods in a unified environment in Section V-B. Note that we only list published works for reference. Following the performance and efficiency comparisons, Section V-C revisits the existing methods according the four core designs of lane detection.
@@ V-B Efficiency Comparison
Considering the variability of equipment used in experiments with existing methods, it is challenging to conduct a relatively fair comparison in terms of execution efficiency. Therefore, we retest representative methods in a unified environment. Table VI shows the work efficiency of these methods. The representative open-source methods are reevaluated according to their settings on CULane or OpenLane dataset. To ensure fairness, only the inference speed of the model is tested to report the frames per second (FPS). The backbone, input size, model’s output, and possible post-processing (whether the model’s output reflects a vectorized representation of each unique lane instance) of each method are also described. All tests utilize a single Nvidia GeForce RTX 3090 GPU.
@@ V-C Discussion
In the two previous chapter, the overview of existing methods is presented from four aspects: task paradigm, lane modeling, global context supplementation, and perspective effect elimination. Combining performance and efficiency comparisons, we continue to discuss their importance for lane detection, as an empirical recipe provided to readers.
$\bullet$ Task Paradigm. segmentation-based methods achieve Instance-level discrimination and lane positioning in a two-stage approach. The majority of the algorithm’s runtime is occupied by independent instance discrimination processes. This makes them overall less efficient than object detection-based methods which achieve in a one-stage. For object detection based methods, it is worth considering the positive and negative sample matching strategy during network training, which will determine whether NMS is needed for post-processing.
$\bullet$ Lane modeling. In mask-based modeling methods [36, 56, 16], each pixel is classified, which can lead to inaccurate segmentation masks that subsequently hinder vectorized fitting. Thus, achieving optimal performance and efficiency remains challenging. In contrast, keypoints-based modeling, line anchor-based modeling, and curve-based modeling methods learn less points or parameters, directly yielding the vectorized results needed downstream.
@@
Line anchor-based modeling methods [80, 19, 25, 27] leverage the vertical and elongated characteristics of lanes in monocular images to strike a good balance between performance and efficiency. Nonetheless, these methods, which learn the horizontal offsets of equidistant points, are unsuitable for U-shaped or nearly horizontal lanes. This corner case further discussed in subsequent sections.
Curve-based modeling methods [82, 20] exhibit decent efficiency but fall short in terms of competitive performance on 2D lane detection benchmarks. Interestingly, this kind of methods achieve strong results in 3D lane detection [130, 125]. As analyzed by Han et al. [140], this discrepancy is due to the ground height influence, which makes fitting irregular lanes challenging in the frontal view (FV). In contrast, these lanes appear smooth in the bird’s-eye view (BEV), where they can be more easily fitted.
@@
Early 3D lane detection methods [22, 23], which still assume a flat ground, leverage IPM to construct BEV features. Some later approaches [24, 26, 125] improve on this by incorporating learnable methods, leading to enhanced performance. Alternatively, other methods [119, 25, 27] avoid BEV feature construction entirely, modeling 3D lanes directly and employing a 3D-to-2D forward projection to circumvent the inherent errors introduced by IPM. It should be pointed out that some methods employ Transformers to achieve view transformation of BEV features [24] or enable interaction between 3D lanes and FV features [27], resulting in improved performance. Nonetheless, the hardware deployment of advanced operators, such as deformable attention [99], remains an problem in need of optimization.
@@ VI-A Multi-task Perception
In autonomous driving, multiple perception tasks often need to be processed synchronously, in real-time, and in parallel. A shared backbone can largely save computation cost and improve efficiency. Thus, leverage a unified framework to conduct multiple perception tasks simultaneously is gradually become a research hotspot. Early works [142, 143, 159, 160, 161, 162] connect multiple specific task heads after the feature extractor to simultaneously complete three tasks on BDD100K [38] dataset: object detection, drivable area segmentation, and lane detection. These methods achieve significant results in each task, which benefit from the powerful and efficient encoder and carefully designed multitasking learning strategy. However, the GT of lanes in the BDD100K are only annotated at the semantic-level, and usually only binary segmentation methods can be used. For lanes, further post-processing is needed to distinguish instances. Recent research mainly focuse on multi-task 3D perception using multiple cameras as sensors. PETRv2 [144] design detection query, segmentation query and lane query to support 3D object detection, BEV segmentation, and 3D lane detection simultaneously. Li et al. [145] propose a unified representation method for multiple perception tasks. They represent 3D objects and 3D lanes as a kind of 3D vector field, which allows they to leverage a single head unified model to achieve multi-task perception.
@@ VI-B Video Lane Detection
As mentioned in the previous chapters, current works attempt to supplement more global information to better detect lanes with unclear visual clues. However, these methods rely on detectors that use single images. In autonomous driving systems, video frames are captured continuously. Therefore, the correlation between frames can be used to more reliably detect obscured lanes in the current frame. For 2D lane detection, Zou et al. [163] and Zhang et al. [164] use recursive neural networks to fuse the features of the current frame with those of several past frames. Zhang et al. [41] aggregate the features of the current frame and multiple past frames based on Transformer. Tabelini et al. [165] extract lane features from video frames using LaneATT [80] and combine these features. Wang et al. [166] utilize spatiotemporal information from adjacent video frames by extending the feature aggregation module in RESA [16]. Jin et al. put RVLD [46], which include an intra frame lane detector to locate lanes in stationary frames and a predictive lane detector to use information from the previous frame for lane detection in the current frame. OMR [152], developed by the same team, employs vehicle masks occupying lanes to interact with historical frames, further improving the accuracy of lane detection in the current frame. For 3D lane detection, STLane3D [167] proposes a multi-frame pre-alignment layer under the BEV space, which uniformly projects features from different frames onto the same ROI region. Anchor3DLane-T [25] incorporates temporal information by projecting the 3D anchors of the current frame onto previous frames to sample features. CurveFormer++ [168] design a temporary Curve Cross Attention module based on CurveFormer [119], which can selectively utilize historical curve query and keypoints to propagate historical information frame by frame.
@@ VI-C Online HD Map Construction
HD maps are an essential module for autonomous driving. Although the traditional offline method of building HD maps can generate accurate map information and is adopted by many autonomous driving companies, it requires a lot of manual annotation cost. As an alternative, Some work try to design a novel HD map learning framework that makes use of on-car sensors and computation to estimate vectorized local semantic maps.
@@
From a process perspective, they typically follow the general pipeline of BEV perception tasks [117], taking multi-camera images as input, extracting image features using a 2D encoder, then obtaining BEV features through a view transform module, and finally outputting various map elements from the BEV perspective through a specific map element decoder. Due to the increase in the number of sensors and the fact that the map elements to be detected include but are not limited to lanes, pedestrian crossings, lane separations, lane boundarys, the task is more challenging than monocular lane detection. Similarly, the key to this type of work is how to model map elements with different shapes, such as lines and polygons, into a set of values that can be learned through neural networks.
@@
HDMapNet [158] adopts a basic bottom-up segmentation approach to perform semantic segmentation on all map elements, and then combines instance embedding and post-processing clustering to obtain each map element instance. This rasterization result still requires post-processing for downstream use, therefore, subsequent work attempts to predict vectorized maps end-to-end. BeMapNet [169] model map elements as segmented Bézier curves. It detects map elements first and then regresses the detailed points with a piecewise Bézier head. VectorMapNet [170] uses a line to represent all map elements and defines a hierarchical query representation. The points of the map elements are autoregressively output through the transformer decoder. However, it outputs the point set through autoregression, which lead to low efficiency. MapTR [148] design a unique representation method for map elements. It uses lines and polygons with uniform sampling points to represent line and area elements, respectively. Therefore, all map elements are represented as sets with the same number of points and different arrangement orders. Owing to its unified permutation-equivalent modeling approach and hierarchical query design, MapTR achieves advanced performance and efficiency on the nuScenes [146] dataset solely with camera input, providing a solid baseline for follow-up research. Afterwards, MapTRv2 [171] improve self attention and cross attention in the decoder of MapTR, further enhancing both the accuracy and performance. PivotNet [149] proposes an end-to-end framework for representing map elements using pivot points. The purpose is to address the issue of shape information loss caused by using a fixed number and consistent position of points to represent complex map elements in MapTR. HIMap [150] meticulously designs feature extractors for MapTR’s hierarchical query, enabling the model to better learn instance level features. StreamMapNet [172] improves MapTR in terms of timing. It overlays information from all historical frames together and implements a memory mechanism using recurrent late embedding. Then MapTracker [151] formalizes the online HD map construction as a tracking task and uses the history of memory latents to ensure consistency in reconstruction over time.
@@ VI-D Lane Topology Reasoning
The early topology reasoning is mainly focused on lane-lane topology, i.e., detecting centerline to construct lane graph. Extraction of the online lane topology structure is first proposed by STSU [173], which extended the task to include centreline extraction and lane connectivity relationships. TopoRoad [174] use a set of directed lane curves and their interactions to represent road topology. Can et al. [175] provide additional supervision of the relationship by considering the centerlines as cluster centers to assign objects. LaneGAP [156] utilizes the shortest path algorithm in graph theory to transform lane topology into a series of overlapping paths, and directly obtains information about these complete paths through end-to-end learning. CenterLineDet [176] regard centerlines as vertices and design a graph model to update centerline topology. Recently, lane-traffic topologys is additionally introduced by the OpenLaneV2 [153] dataset, further improving the understanding of scene structure. Aiming at a complete and diverse driving scene graph, TopoNet  [177] explicitly models the connectivity of centerlines within the network and incorporates traffic elements into the task. TopoMLP  [154] leverage position embedding [138] to enhance topology modeling. LaneSegNet [155] propose a unified representation that integrates lanes and centerlines, and introduced a lane attention mechanism to facilitate the learning of topological relationships between centerlines and lanes. To improve lane topology inference, TopoLogic [157] introduces an efficient post-processing that integrates the geometric distance between centerline endpoints and the semantic similarity of lane queries within a high-dimensional space.

### v2

@@ IV-B2 Learnable View Transformation
To reduce the inherent errors caused by this rigid transformation, some researchers adopt a more flexible approach. They use neural networks to learn the transformation process from FV to BEV features. PersFormer [24] leverages deformable attention [99] to learn the spatial transformation from FV to BEV. It references the coordinate transformation matrix of IPM to generate BEV feature representations, focusing on relevant regions in the FV features. In the lane decoding stage, PersFormer adopts a unified 2D/3D line anchor design, achieving unified 2D and 3D lane detection. BEV-LaneDet [26] integrates the MLP based view transformation method VPN [120] into FPN [131] to obtain BEV features. It constructs a virtual camera module to project all images onto a standard virtual camera view, ensuring consistent image distribution. For lane modeling, BEV-LaneDet models lanes as keypoints. It adopts an embedding-clustering-based instance segmentation method [65] and refers to YOLO [96] to divide the BEV plane into grids, predicting the offset of each grid’s center point relative to GT. Due to its concise architecture, it is well-suited for deployment. Yao et al. [121] add the coarse-to-fine mechanism [19] based on PersFormer [24]. They fuse the local and global information referring to the coordinates of sparse points and jointly refine the global and local structures of lanes. Chen et al. [132] decompose the cross-attention between FV and BEV features into separate cross-attentions: one between FV and lane features, and another between BEV and lane features. Dynamic kernels are then used to convolve FV and BEV feature maps, generating 2D and 3D lane keypoint offset maps. GroupLane [133] uses the depth-estimation-based view transformation method LSS [126] to obtain BEV features. It models BEV lanes as grids [68] and establishes vertical and horizontal group heads to identify horizontal and vertical lanes, respectively. LaneCPP [125] also completes BEV transformation based on LSS. It models lanes as B-spline curves and uses prior knowledge of road geometry to enhance view transformation and lane prediction.
@@ IV-C2 Derictly Modeling 3D Lanes
The 3D object detection methods based on sparse BEV representations, such as DETR3D [137] and PETR [138], guide this approach. CurveFormer [119] constructs curve queries by modeling 3D lanes as 3D line anchors to provide explicit positional priors [139]. It designs a curve cross-attention mechanism to predict polynomial parameters for 3D lanes. Inspired by the line anchor feature pooling mechanism in LaneATT [80], Anchor3DLane [25] utilizes camera intrinsics and extrinsics to accurately project 3D line anchor points onto FV features. This facilitates anchor feature sampling, allowing the network to predict 3D coordinates and lane classification results based on the sampled anchor features. LATR [27] decomposes 3D line anchors into dynamically generated point-level and lane-level queries. It uses dynamic 3D ground position embeddings to interact with FV features, updating the lane query to bridge 3D space and 2D images. Dong et al. [124] model lanes as 3D Bézier curves and predict the curve control points via Transformer. Han et al. [140] expressed the 3D lane as a polynomial in 3D space. They design two Transformer structures to learn the 2D polynomial and height of the X-O-Z plane and project the resulting 3D lane onto FV for supervised alignment. PVALane [123] generates sparser 3D line anchors than Anchor3DLane by first predicting 2D lanes in FV. It also introduces a module to align sampled FV and BEV features for more accurate 3D lane detection.
@@ V Benchmark Results
This section reports the performance of representative lane detection methods on commonly used public datasets. For each reviewed area, the most widely used datasets are selected for benchmarking in Section V-A. Because of the high efficiency requirement for lane detection, speed tests are also conducted on representative open-source lane detection methods in a unified environment in Section V-B. Note that we only list published works for reference. Following the performance and efficiency comparisons, Section V-C revisits the existing methods according to the four core designs of lane detection.
@@ V-B Efficiency Comparison
Considering the variability of equipment used in experiments with existing methods, it is challenging to conduct a relatively fair comparison in terms of execution efficiency. Therefore, we retest representative methods in a unified environment. Table VI shows the work efficiency of these methods. The representative open-source methods are reevaluated according to their settings on the CULane or OpenLane dataset. To ensure fairness, only the inference speed of the model is tested to report the frames per second (FPS). The backbone, input size, model’s output, and possible post-processing (whether the model’s output reflects a vectorized representation of each unique lane instance) of each method are also described. All tests utilize a single Nvidia GeForce RTX 3090 GPU.
@@ V-C Discussion
In the two previous chapters, the overview of existing methods is presented from four aspects: task paradigm, lane modeling, global context supplementation, and perspective effect elimination. Combining performance and efficiency comparisons, we continue to discuss their importance for lane detection, as an empirical recipe provided to readers.
$\bullet$ Task Paradigm. segmentation-based methods achieve Instance-level discrimination and lane positioning in a two-stage approach. The majority of the algorithm’s runtime is occupied by independent instance discrimination processes. This makes them overall less efficient than object detection-based methods which are achieved in a one-stage. For object detection-based methods, it is worth considering the positive and negative sample matching strategy during network training, which will determine whether NMS is needed for post-processing.
$\bullet$ Lane modeling. In mask-based modeling methods [36, 56, 16], each pixel is classified, which can lead to inaccurate segmentation masks that subsequently hinder vectorized fitting. Thus, achieving optimal performance and efficiency remains challenging. In contrast, keypoints-based modeling, line anchor-based modeling, and curve-based modeling methods learn fewer points or parameters, directly yielding the vectorized results needed downstream.
@@
Line anchor-based modeling methods [80, 19, 25, 27] leverage the vertical and elongated characteristics of lanes in monocular images to strike a good balance between performance and efficiency. Nonetheless, these methods, which learn the horizontal offsets of equidistant points, are unsuitable for U-shaped or nearly horizontal lanes. This corner case is further discussed in subsequent sections.
Curve-based modeling methods [82, 20] exhibit decent efficiency but fall short in terms of competitive performance on 2D lane detection benchmarks. Interestingly, this kind of method achieves strong results in 3D lane detection [130, 125]. As analyzed by Han et al. [140], this discrepancy is due to the ground height influence, which makes fitting irregular lanes challenging in the frontal view (FV). In contrast, these lanes appear smooth in the bird’s-eye view (BEV), where they can be more easily fitted.
@@
Early 3D lane detection methods [22, 23], which still assume a flat ground, leverage IPM to construct BEV features. Some later approaches [24, 26, 125] improve on this by incorporating learnable methods, leading to enhanced performance. Alternatively, other methods [119, 25, 27] avoid BEV feature construction entirely, modeling 3D lanes directly and employing a 3D-to-2D forward projection to circumvent the inherent errors introduced by IPM. It should be pointed out that some methods employ Transformers to achieve view transformation of BEV features [24] or enable interaction between 3D lanes and FV features [27], resulting in improved performance. Nonetheless, the hardware deployment of advanced operators, such as deformable attention [99], remains a problem in need of optimization.
@@ VI-A Multi-task Perception
In autonomous driving, multiple perception tasks often need to be processed synchronously, in real-time, and parallel. A shared backbone can largely save computation costs and improve efficiency. Thus, leveraging a unified framework to conduct multiple perception tasks simultaneously gradually becomes a research hotspot. Early works [142, 143, 159, 160, 161, 162] connect multiple specific task heads after the feature extractor to simultaneously complete three tasks on BDD100K [38] dataset: object detection, drivable area segmentation, and lane detection. These methods achieve significant results in each task, which benefit from the powerful and efficient encoder and carefully designed multitasking learning strategy. However, the GT of lanes in the BDD100K are only annotated at the semantic-level, and usually only binary segmentation methods can be used. For lanes, further post-processing is needed to distinguish instances. Recent research mainly focuses on multi-task 3D perception using multiple cameras as sensors. PETRv2 [144] design detection query, segmentation query, and lane query to support 3D object detection, BEV segmentation, and 3D lane detection simultaneously. Li et al. [145] propose a unified representation method for multiple perception tasks. They represent 3D objects and 3D lanes as a kind of 3D vector field, which allows them to leverage a single-head unified model to achieve multi-task perception.
@@ VI-B Video Lane Detection
As mentioned in the previous chapters, current works attempt to supplement more global information to better detect lanes with unclear visual clues. However, these methods rely on detectors that use single images. In autonomous driving systems, video frames are captured continuously. Therefore, the correlation between frames can be used to more reliably detect obscure lanes in the current frame. For 2D lane detection, Zou et al. [163] and Zhang et al. [164] use recursive neural networks to fuse the features of the current frame with those of several past frames. Zhang et al. [41] aggregate the features of the current frame and multiple past frames based on Transformer. Tabelini et al. [165] extract lane features from video frames using LaneATT [80] and combine these features. Wang et al. [166] utilize spatiotemporal information from adjacent video frames by extending the feature aggregation module in RESA [16]. Jin et al. put RVLD [46], which includes an intra-frame lane detector to locate lanes in stationary frames and a predictive lane detector to use information from the previous frame for lane detection in the current frame. OMR [152], developed by the same team, employs vehicle masks occupying lanes to interact with historical frames, further improving the accuracy of lane detection in the current frame. For 3D lane detection, STLane3D [167] proposes a multi-frame pre-alignment layer under the BEV space, which uniformly projects features from different frames onto the same ROI region. Anchor3DLane-T [25] incorporates temporal information by projecting the 3D anchors of the current frame onto previous frames to sample features. CurveFormer++ [168] design a temporary Curve Cross Attention module based on CurveFormer [119], which can selectively utilize historical curve query and keypoints to propagate historical information frame by frame.
@@ VI-C Online HD Map Construction
HD maps are an essential module for autonomous driving. Although the traditional offline method of building HD maps can generate accurate map information and is adopted by many autonomous driving companies, it requires a lot of manual annotation costs. As an alternative, Some work try to design a novel HD map learning framework that makes use of on-car sensors and computation to estimate vectorized local semantic maps.
@@
From a process perspective, they typically follow the general pipeline of BEV perception tasks [117], taking multi-camera images as input, extracting image features using a 2D encoder, then obtaining BEV features through a view transform module, and finally outputting various map elements from the BEV perspective through a specific map element decoder. Due to the increase in the number of sensors and the fact that the map elements to be detected include but are not limited to lanes, pedestrian crossings, lane separations, and lane boundaries, the task is more challenging than monocular lane detection. Similarly, the key to this type of work is how to model map elements with different shapes, such as lines and polygons, into a set of values that can be learned through neural networks.
@@
HDMapNet [158] adopts a basic bottom-up segmentation approach to perform semantic segmentation on all map elements. Then it combines instance embedding and post-processing clustering to obtain each map element instance. This rasterization result still requires post-processing for downstream use, therefore, subsequent work attempts to predict vectorized maps end-to-end. BeMapNet [169] model map elements as segmented Bézier curves. It detects map elements first and then regresses the detailed points with a piecewise Bézier head. VectorMapNet [170] uses a line to represent all map elements and defines a hierarchical query representation. The points of the map elements are autoregressively output through the transformer decoder. However, it outputs the point set through autoregression, which leads to low efficiency. MapTR [148] design a unique representation method for map elements. It uses lines and polygons with uniform sampling points to represent line and area elements, respectively. Therefore, all map elements are represented as sets with the same number of points and different arrangement orders. Owing to its unified permutation-equivalent modeling approach and hierarchical query design, MapTR achieves advanced performance and efficiency on the nuScenes [146] dataset solely with camera input, providing a solid baseline for follow-up research. Afterward, MapTRv2 [171] improves self-attention and cross-attention in the decoder of MapTR, further enhancing both the accuracy and performance. PivotNet [149] proposes an end-to-end framework for representing map elements using pivot points. The purpose is to address the issue of shape information loss caused by using a fixed number and consistent position of points to represent complex map elements in MapTR. HIMap [150] meticulously designs feature extractors for MapTR’s hierarchical query, enabling the model to better learn instance-level features. StreamMapNet [172] improves MapTR in terms of timing. It overlays information from all historical frames together and implements a memory mechanism using recurrent late embedding. Then MapTracker [151] formalizes the online HD map construction as a tracking task and uses the history of memory latents to ensure consistency in reconstruction over time.
@@ VI-D Lane Topology Reasoning
The early topology reasoning is mainly focused on lane-lane topology, i.e., detecting the centerlines to construct a lane graph. Extraction of the online lane topology structure is first proposed by STSU [173], which extended the task to include centreline extraction and lane connectivity relationships. TopoRoad [174] uses a set of directed lane curves and their interactions to represent road topology. Can et al. [175] provide additional supervision of the relationship by considering the centerlines as cluster centers to assign objects. LaneGAP [156] utilizes the shortest path algorithm in graph theory to transform lane topology into a series of overlapping paths and directly obtains information about these complete paths through end-to-end learning. CenterLineDet [176] regard centerlines as vertices and design a graph model to update centerline topology. Recently, lane-traffic topology is additionally introduced by the OpenLaneV2 [153] dataset, further improving the understanding of scene structure. Aiming at a complete and diverse driving scene graph, TopoNet  [177] explicitly models the connectivity of centerlines within the network and incorporates traffic elements into the task. TopoMLP  [154] leverage position embedding [138] to enhance topology modeling. LaneSegNet [155] proposes a unified representation for integrating lanes and centerlines. It introduces a lane attention mechanism to facilitate the learning of topological relationships between centerlines and lanes. To improve lane topology inference, TopoLogic [157] introduces an efficient post-processing that integrates the geometric distance between centerline endpoints and the semantic similarity of lane queries within a high-dimensional space.